CH

3OH, is 35.20 kJ/mol. Its boiling

point is 64.6°C


1
Expert's answer
2021-06-21T05:31:44-0400

Solution


calculating the entropy methanol

"∆S_{CH_3OH}=\frac{∆_{vap}}{T}"

T in Kelvin=64.6+273=336.6K


"∆S_{CH_3OH}= \frac{35.20\times1000}{336.6}"


Entropy of methanol

=104.58J/mol.K